OpenStudy (anonymous):

Algebra 2 help, please?

OpenStudy (anonymous):

What is the question that you need help with?

OpenStudy (anonymous):

whats your question?

OpenStudy (anonymous):

\[\frac{ 5x^2+9x-2 }{ x^2+12x+20 }\times \frac{ x^2+17x+70 }{ 15x-3 }\]

OpenStudy (anonymous):

Just to simplify that

OpenStudy (anonymous):

Sorry, I'm to lazy to work that WHOLE thing out :(

OpenStudy (anonymous):

Tip though, multiply the top of the first by x^2, then by 17x, and then 70. Do the same thing with the bottom, and divide the two (First one on the top)

OpenStudy (anonymous):

Ok, thank you